GDP per capita in Latvia?

GDP per capita in the country?

GDP growth in Latvia%?

Growth of GDP per capita in Latvia%?

GDP growth in the country%?

Growth of GDP per capita in the country%?

Number of years when Latvia achieves the same level (based on GDP growth)?

Number of years when Latvia achieves the same level ( based on growth of GDP per capita)?

The number must be written as% but without the simbol%
Answer #1

As per data received from World Bank Official websjte which has released macro data for Lativia in year 2017.

Gdp Per Capita in 2017= 16,800 dollars

Gdp in 2017 = 30.2 billion dollars

GDP per capita in 2016= 15,900 dollars

Gdp in 2016 =27.5 billion dollars

GDP growth over 2016 = 30.2-27.5/27.5= 9.8 percent

GDP growth per capita= 16,800 - 15900/15900= 5.6 percent

Average GDP per capita in Latvia and nearby countries= 18500 dollars

Average GDP in nearby countries = 32. 4 billion dollars

Assuming grwoth of 9.8 percent from 2017 onwards, Latvia shall cross average GDp in nearby countries within 1 year.

And shall cross GDP per capita in nearby countries within 2 years approximately.